OpenStudy (anonymous):

Which of the pairs of ratios are equivalent? a.6 : 7, 30 : 36 b.18 to 24, 24 to 30 c.18 to 24, 3 to 4 d.6/7,42/56 d.

OpenStudy (anonymous):

@chosenmatt

OpenStudy (aripotta):

does 6/7 = 30/36?

OpenStudy (anonymous):

so a

OpenStudy (aripotta):

no...but 6/7 doesn't equal 30/36. i want you to go through each answer and ask yourself if they're equal or not

OpenStudy (chosenmatt):

@AriPotta is right good job :)

OpenStudy (aripotta):

6/7, supposedly, would have to be multiplied by 5/5, right? but 6/7 x 5/5 = 30/35, which is not 30/36

OpenStudy (aripotta):

so a is not correct

OpenStudy (aripotta):

try b yourself. is b right?